A rock is thrown horizontally from the top of a cliff 88m high , with a horizontal speed of 24m/s with what velocity does the rock hit and what is the direction of velocity


1
Expert's answer
2022-05-11T13:52:34-0400

"v_x=24~\frac ms,"

"v_y=\sqrt{2gh}=48~\frac ms,"

"v=\sqrt{v_x^2+v_y^2}=53.7~\frac ms,"

"\alpha=-\arctan\frac{v_y}{v_x}=-63.4°."